# Revisiting Reset Mechanisms in Spiking Neural Networks for Sequential Modeling: Specialized Discretization for Binary Activated RNN

### v2

@@ 1 Declarative Network Perspective of Two-Phase TTFS-based Spiking Neural Networks
This subsection is not directly related to the main content of this article but serves as an example of the quantized activation value perspective in spiking neural networks.
Two-stage TTFS-based spiking neural networks refer to networks where each layer operates in a separate time phase, ensuring that each neuron in each layer fires at most one spike. However, this requires longer inference time steps because each layer needs its own dedicated phase for computation. Previously discussed rate-coded spiking neural networks can be easily understood through the lens of frequency as quantized activations in traditional neural networks. Here, we further explain the declarative network nature of temporally coded spiking neural networks—specifically, two-phase TTFS-based spiking neural networks,which have achieved strong performance, helping readers better understand the relationship between spiking neural networks and quantized neural networks. For readers interested in the specifics of two-stage TTFS-based spiking neural networks, please refer to literature such as [33, 34, 35]. These models typically consist of two phases: the first phase receives spike inputs, often using a linear increase in membrane potential with a slope determined by parameter W, followed by the second phase, the spike generation phase, where the neuron fires its only spike once the linearly increasing membrane potential reaches the threshold. We will not delve into further details here. Instead, we aim to highlight that this second phase can be viewed as a step-by-step search for the solution to an equation, thus making it interpretable as a declarative network with quantized activations.
@@
Spiking neural networks that emphasize temporal information and employ temporal encoding can also be viewed as a form of declarative network. These models focus on spike timing, where the firing times of neurons in each layer encode the layer’s information. The key question then becomes: when should each neuron in a layer fire? Thus, temporally encoded spiking neural networks can be interpreted as solving an equation to compute the spike timing of neurons in each layer as its output. Specifically, they solve the equation $f(x,y,\theta)=0$, where $y$ is the variable to be solved (i.e., the spike timing), $x$ represents the input spike timing, and $\theta$ denotes the learnable parameters.

### v4

@@ 1 Declarative Network Perspective of Two-Phase TTFS-based Spiking Neural Networks
This subsection is not directly related to the main content of this article but serves as an example of the quantized activation value perspective in spiking neural networks. Additionally, my previous paper[49] had some shortcomings in terms of rigor and detail, so I revisit and present its core content in Appendices 1 and 2 through a more detailed analysis. Appendix 1 shows that if a spiking neural network employs separate timesteps for each layer, such as in the two-phase Time-to-First-Spike (TTFS) scheme, it can, in essence, be functionally equivalent to many quantized declarative artificial neural networks. Appendix 2 offers a straightforward, distribution-based perspective to understand the fundamental nature of information encoding in spike trains.
Two-stage TTFS-based spiking neural networks refer to networks where each layer operates in a separate time phase, ensuring that each neuron in each layer fires at most one spike. However, this requires longer inference time steps because each layer needs its own dedicated phase for computation. Previously discussed rate-coded spiking neural networks can be easily understood through the lens of frequency as quantized activations in traditional neural networks. Here, we further explain the declarative network nature of temporally coded spiking neural networks—specifically, two-phase TTFS-based spiking neural networks,which have achieved strong performance, helping readers better understand the relationship between spiking neural networks and quantized neural networks. For readers interested in the specifics of two-stage TTFS-based spiking neural networks, please refer to literature such as [33, 34, 35]. These models typically consist of two phases: the first phase receives spike inputs, often using a linear increase in membrane potential with a slope determined by parameter W, followed by the second phase, the spike generation phase, where the neuron fires its only spike once the linearly increasing membrane potential reaches the threshold. We will not delve into further details here. Instead, we aim to highlight that this second phase can be viewed as a step-by-step search for the solution to an equation, thus making it interpretable as a declarative network[37] with quantized activations.
@@
Spiking neural networks that emphasize temporal information and employ temporal encoding can also be viewed as a form of declarative network[51]. These models focus on spike timing, where the firing times of neurons in each layer encode the layer’s information. The key question then becomes: when should each neuron in a layer fire? Thus, temporally encoded spiking neural networks can be interpreted as solving an equation to compute the spike timing of neurons in each layer as its output. Specifically, they solve the equation $f(x,y,\theta)=0$, where $y$ is the variable to be solved (i.e., the spike timing), $x$ represents the input spike timing, and $\theta$ denotes the learnable parameters.